OpenStudy (anonymous):

An integer is chosen at random from the first 40 positive integers. What is the probability that the integer chosen is divisible by 6 or 8?

OpenStudy (anonymous):

im thinking 11/40 but i read somewhere else its 9/40. Could you please explain how you get your answer

OpenStudy (anonymous):

how many of the 1st 40 integers are divisible by 6? how many by 8? how many by both?

OpenStudy (anonymous):

P(A or B) = P(A) + P(B) - P(A and B)

OpenStudy (anonymous):

6 are divisible by 6. 5 are divisible by 8 correct?

OpenStudy (anonymous):

yes and how many are divisible by both 6 and 8?

OpenStudy (anonymous):

umm. is there an easy way to find out?

OpenStudy (anonymous):

every 4th multiple of 6 should also be divible by 8 or (to think of it in a different way) every 3rd multiple of 8 should be divisible by 6. why? 6*4k = 2*4*3k = 8*3k => multiple of 8 8*3k = 4*2*3k = 6*4k => multiple of 6

OpenStudy (anonymous):

so only one is divisible by 6 and 8.

OpenStudy (anonymous):

so the answer would be 10/40 or 1/4.... correct?

OpenStudy (anonymous):

yep! good job!

OpenStudy (anonymous):

ok thanks! Im going to ask another that is more complicated. if you could help i would appreciate it:)

OpenStudy (anonymous):

just @ me in your post and i will help